A 60.0 g sample containing mixed BaO and CaO are transformed into mixed sulfates weighing 10.34 grams. Calculate the percentage BaO and CaO in the sample.
Answer #1

Bao= x* 153.3
Cao= y*56.0779

x*153.3g + y*56.0774=10.34g

x*233.38g + y*136.14=60.00g

x=126.744 molBao

y= 185.8 molCao